15.1 Using the Straight Line (SL) depreciation method, what is the (a) annual depreciation charge, Dt=?, (b) the total accumulated depreciation charge, D4=?, at the end of year 4, and (c) the Book Value, BV=?, at the end of year 4, for an automobile with a five-year useful life that has a first cost of $20,084 and an $8,601 salvage value?
Answer #1

P = 20084, S = 8601, N = 5

a) Depreciation per year = (P-S)/N = (20084 - 8601)/5 = 11483/5 = 2296.60

b) Total depreciation after 4 yrs = 4 * 2296.60 = 9186.40

c) Book value after 4 yrs = 20084 - 9186.60 = 10897.60